OpenStudy (anonymous):

Which function below is the inverse of f(x) = x2 − 16?

OpenStudy (anonymous):

Do you know what an inverse is?

OpenStudy (anonymous):

The Opp Basically ..

OpenStudy (anonymous):

Kind of. With an inverse, you replace x with y and solve. So x=y^2+16 \[x=y^2+16\] \[x-16=y^2\] \[\sqrt{x-16}=\sqrt{y^2}\] \[y=\sqrt{x-16}\]

OpenStudy (anonymous):

thank you !!

OpenStudy (anonymous):

i dont think this is correct ??

OpenStudy (anonymous):

x squared over 16 ±4square root of x ±square root of the quantity x plus 16 1 over quantity x squared minus 16

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=inverse+of+x^2%2B16&dataset= I just Wolfram Alpha'd it and it's correct.

OpenStudy (anonymous):

i have no idea what Wolfram Alpha is but okayy ( I dont see an answer like it

OpenStudy (anonymous):

Wolfram Alpha is basically a website that can give you an answer to almost anything. But I think your teacher is wrong. Or you put in the wrong question.

OpenStudy (anonymous):

its a worksheet i didnt know how to do

OpenStudy (anonymous):

@iPwnBunnies Did I do something wrong here or?

OpenStudy (anonymous):

no im just not sure if your web site is correct but thank you anyway

OpenStudy (anonymous):

I'm 99.9999% sure that my website is correct.

OpenStudy (anonymous):

@jim_thompson5910 Am I wrong? .-.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Or @mathmale can you check this out since it looks like everyone else is afk?

jimthompson5910 (jim_thompson5910):

f(x) = x^2 − 16 is the original function, not f(x) = x^2 + 16

jimthompson5910 (jim_thompson5910):

also, remember that \[\Large x^2 = k \implies x = \pm \sqrt{k}\]

jimthompson5910 (jim_thompson5910):

even though technically the plus/minus doesn't make it a function (so they contradict themselves here)

OpenStudy (anonymous):

Ah so yes, I was wrong. Your answer would be +/- sqrt(x+16).

jimthompson5910 (jim_thompson5910):

what's frustrating is that it should be \[\Large y = \sqrt{x+16}\] assuming x > 0 in the original function. This is because they ask for an inverse FUNCTION. However, the answer choice \[\Large y = \pm \sqrt{x+16}\] is NOT a function because of the plus/minus. So they are contradicting themselves.

OpenStudy (anonymous):

Well regular functions have the vertical line test to check if it's a function, don't they have the horizontal line test for inverse functions?

jimthompson5910 (jim_thompson5910):

yes if the function fails the horizontal line test, then it doesn't have an inverse

jimthompson5910 (jim_thompson5910):

restricting the domain changes that though

jimthompson5910 (jim_thompson5910):

so that's why trig functions (which are periodic and aren't one-to-one in nature) have inverse functions

OpenStudy (anonymous):

Hmm. Doesn't that mean a parabola (say, x^2) wouldn't have have an inverse? Yet it does, sqrt (x)